OpenStudy (anonymous):

integral from 1 to 4 of: 1/(sqrtx)(sqrtx+1)^3

OpenStudy (anonymous):

\[\frac{ 1 }{ (\sqrt{x})(\sqrt{x+1)}^{3} }\]

OpenStudy (anonymous):

\[\int\limits_{1}^{4} \frac{ 1 }{ \sqrt{x} (\sqrt{x+1))^{3}}}\] Is this how you write it?

OpenStudy (anonymous):

Yes, only sqrtx+1 is raised to the third

OpenStudy (anonymous):

I know to use u-substitution but other than that I'm lost on how to solve this

OpenStudy (anonymous):

so far